OpenStudy (anonymous):

Combine like terms: x + 5 x minus 2 x + 3 minus 4. A. 4 x minus 1 B. 3 x plus 1 C. 3 x minus 1 D. negative 10 x minus 12

OpenStudy (texaschic101):

x + 5x - 2x + 3 - 4 combine like terms (x is the same as 1x) so we have 1x + 5x - 2x = 6x - 2x = 4x and then we have 3 - 4 = -1 so put them together and we have : 4x - 1 any questions ?

OpenStudy (anonymous):

thank you so much

OpenStudy (texaschic101):

no problem :)